# Quantum Gravity, Hydrodynamics and Emergent Cosmology: A Collection of Perspectives

### v1

@@ Acknowledgements
The work of Jibril Ben Achour is supported by the Sir John Templeton foundation and the MCQST through the seed funding program. Flaminia Giacomini acknowledges support from the Swiss National Science Foundation via the Ambizione Grant PZ00P2-208885, the ETH Zurich Quantum Center, and from the John Templeton Foundation, as part of the ‘The Quantum Information Structure of Spacetime, Second Phase (QISS 2)’ Project. Christophe Goeller acknowledges support by the Alexander von Humboldt Foundation and the MCQST for its support through the seed funding program. Tobias Haas acknowledges support from the European Union under project ShoQC within the ERA-NET Cofund in Quantum Technologies (QuantERA) program and from the F.R.S.- FNRS under project CHEQS within the Excellence of Science (EOS) program. Luca Marchetti acknowledges support from the Atlantic Association for Research in Mathematical Sciences. Daniele Oriti acknowledges financial support from the ATRAE programme of the Spanish Government, through the grant PR28/23 ATR2023-145735. Daniele Oriti and Andreas Pithis acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) research grants OR432/3-1 and OR432/4-1 and the John-Templeton Foundation via research grant 6242. Antonio D. Pereira acknowledges CNPq under the grant PQ-2 (312211/2022-8), FAPERJ under the “Jovem Cientista do Nosso Estado” program (E26/202.800/2019 and E-26/205.924/2022) and NWO under the VENI Grant (VI.Veni.192.109) for financial support. Andreas Pithis is grateful for the generous financial support by the MCQST via the seed funding through the DFG under Germany’s Excellence Strategy – EXC-2111 – 390814868 and in particular acknowledges funding by the DFG under the author’s project number 527121685 as a Principal Investigator. Sebastian Steinhaus gratefully acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) project number 422809950 and Johannes Thürigen is funded by DFG grant number 418838388 and Germany’s Excellence Strategy EXC 2044–390685587, Mathematics Münster: Dynamics–Geometry–Structure.

### v2

@@ Acknowledgements
The work of Jibril Ben Achour is supported by the Sir John Templeton foundation and the MCQST through the seed funding program. Flaminia Giacomini acknowledges support from the Swiss National Science Foundation via the Ambizione Grant PZ00P2-208885, the ETH Zurich Quantum Center, and from the John Templeton Foundation, as part of the ‘The Quantum Information Structure of Spacetime, Second Phase (QISS 2)’ Project. Christophe Goeller acknowledges support by the Alexander von Humboldt Foundation and the MCQST for its support through the seed funding program. Tobias Haas acknowledges support from the European Union under project ShoQC within the ERA-NET Cofund in Quantum Technologies (QuantERA) program and from the F.R.S.- FNRS under project CHEQS within the Excellence of Science (EOS) program. Philipp Höhn acknowledges support of the ID# 62312 grant from the John Templeton Foundation, as part of the project ‘The Quantum Information Structure of Spacetime’ (QISS). The opinions expressed in this publication are those of the Philipp Höhn and do not necessarily reflect the views of the John Templeton Foundation. Luca Marchetti acknowledges support from the Atlantic Association for Research in Mathematical Sciences. Daniele Oriti acknowledges financial support from the ATRAE program of the Spanish Government, through the grant PR28/23 ATR2023-145735. Daniele Oriti and Andreas Pithis acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) research grants OR432/3-1 and OR432/4-1 and the John-Templeton Foundation via research grant 6242. Antonio D. Pereira acknowledges CNPq under the grant PQ-2 (312211/2022-8), FAPERJ under the “Jovem Cientista do Nosso Estado” program (E26/202.800/2019 and E-26/205.924/2022) and NWO under the VENI Grant (VI.Veni.192.109) for financial support. Andreas Pithis is grateful for the generous financial support by the MCQST via the seed funding through the DFG under Germany’s Excellence Strategy – EXC-2111 – 390814868 and in particular acknowledges funding by the DFG under the author’s project number 527121685 as a Principal Investigator. Sebastian Steinhaus gratefully acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) project number 422809950 and Johannes Thürigen is funded by DFG grant number 418838388 and Germany’s Excellence Strategy EXC 2044–390685587, Mathematics Münster: Dynamics–Geometry–Structure.